OpenStudy (anonymous):

The price of a house is originally listed at $190,000. The owners are having a hard time selling it and decide to reduce the price to $178,600. What is the percentage decrease of the price of the house? a 0.6% b 94% c 6% d 11.4%

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (anonymous):

@Mimi_x3

OpenStudy (clongoria2):

178600 / 190000 = ?

OpenStudy (anonymous):

The decrease in price is the original price minus the new price. Once you know what that number is, you can find out what the percentage of decrease is using a proportion. So ask yourself what is the whole amount and with is the part. \[\frac{ part }{ whole }= \frac{ percentage of decrease }{ 100% }\]